how many grams of O2 will react completely with 35.0 grams of PH3?



1
Expert's answer
2022-02-14T17:03:04-0500

4PH3+8O2=P4O10+6H2O

35g of PH3

MW PH3:34g/mol

mol PH3=

mol =mass/MW

=35/34=1.03

mol ratio PH3:O2=4:8

mol of O2=8/4 ×1.03 =2.06

MW O2=32g/mol

Grams of O2:

gr=mol\MW

=2.06\32

=65.92 grams